### v4

# Automate Strategy Finding with LLM in Quant Investment

Abstract
We present a novel three-stage framework leveraging Large Language Models (LLMs) within a risk-aware multi-agent system for automate strategy finding in quantitative finance. Our approach addresses the brittleness of traditional deep learning models in financial applications by: employing prompt-engineered LLMs to generate executable alpha factor candidates across diverse financial data, implementing multimodal agent-based evaluation that filters factors based on market status, predictive quality while maintaining category balance, and deploying dynamic weight optimization that adapts to market conditions. Experimental results demonstrate the robust performance of the strategy in Chinese & US market regimes compared to established benchmarks. Our work extends LLMs capabilities to quantitative trading, providing a scalable architecture for financial signal extraction and portfolio construction. The overall framework significantly outperforms all benchmarks with 53.17% cumulative return on SSE50  [^1] The SSE 50 Index tracks the performance of the 50 most influential large-cap blue-chip stocks on the Shanghai Stock Exchange. (Jan 2023 to Jan 2024), demonstrating superior risk-adjusted performance and downside protection on the market. Our protocol code are available at https://github.com/kouzhizhuo/Automate-Strategy-Finding-with-LLM-in-Quant-investment.
Automate Strategy Finding with LLM in Quant Investment
Zhizhuo Kou1, Holam Yu2, Junyu Luo3, Jingshu Peng1, Xujia Li1, Chengzhong Liu1, Juntao Dai3, Lei Chen1, 2, Sirui Han1thanks: Corresponding author. thanks: Project leader Yike Guo [^1] footnotemark: 1, 1 The Hong Kong University of Science and Technology 2 The Hong Kong University of Science and Technology (Guangzhou) 3 Peking University zkouaa@connect.ust.hk {siruihan, yikeguo}@ust.hk
1 Introduction
Recent advances in LLMs and multi-agent systems are converging to transform quantitative finance. This synergistic relationship leverages LLMs’ text comprehension and generation capabilities alongside multi-agent frameworks that simulate market dynamics, creating sophisticated approaches to portfolio management Lee et al. (2020).
LLMs have evolved from supporting analytical tools to active participants in financial decision-making Luo et al. (2025c). For example, BloombergGPT demonstrates superior performance in parsing market sentiment and answering domain-specific questions Wu et al. (2023). Research shows that LLMs effectively generate trading actions by contextualizing price trends with news and earnings reports Ding et al. (2023). Concurrently, multi-agent systems offer powerful approaches to portfolio optimization through decentralized interaction. The Multi-Agent Portfolio System demonstrates the portfolio managed by agents have achieved well-diversified returns with improved risk-adjusted performance Lee et al. (2025). These frameworks capture complex market dynamics such as information sharing and strategic arbitrage that single-agent models cannot address Spooner and Savani (2020). The integration of these technologies creates sophisticated financial environments where LLM-enhanced agents demonstrate adaptive behavior. StockAgent exemplifies this approach with LLM-powered agents mimic diverse investor personas responding to market events Zhang et al. (2024a). Hierarchical structures such as FinCon organize agents in manager-analyst relationships, facilitating collaboration through natural language communication Yu et al. (2024). This convergence heralds a future of intelligent, distributed financial decision making that combines data-driven learning with human-like reasoning capabilities for more robust investment strategies Yu et al. (2023b).
Alpha mining is the process of discovering trading signals that generate excess returns in financial markets, we identify three critical challenges in alpha mining: Rigidity of traditional methods that lack adaptability to dynamic markets Tang et al. (2025); Data diversity and integration challenges despite machine learning advances Cui et al. (2021); Yu et al. (2023a); and Adaptation to market variability despite progress in prediction Xu and Cohen (2018) and strategy formulation Chen et al. (2019).
Our LLM-driven framework addresses these limitations through three components: Flexible Alpha Mining employs LLMs to extract, categorize, and filter alpha factors from financial literature, organizing them as momentum, fundamental, or liquidity factors with established independence Xu and Cohen (2018). Multi-agent Multimodal Market Evaluation conducts rigorous backtesting across diverse market conditions, with specialized agents evaluating factor effectiveness from multiple perspectives. Dynamic Strategy Optimization implements a weight gating layer that assigns optimal alpha factor weights based on current market conditions, ensuring adaptive strategy development.
Our methodology synthesizes cutting-edge machine learning techniques with established financial domain knowledge to create a robust interdisciplinary framework for alpha identification and optimization across diverse asset classes. This research is grounded in empirical quantitative investment practices, bridging theoretical advancements with practical applications in portfolio management.
Our main contributions are three-fold: ❶ A novel framework for identifying formulaic alpha factors using LLMs, leveraging their exploratory capabilities to establish an Alpha factory from multimodal information with incremental update functionality; ❷ Introduction of a multi-agent approach to portfolio management for evaluating relationships between market conditions and alpha factors, enabling specialized evaluation under different scenarios; ❸ Integration of advanced techniques from machine learning and finance, representing a significant advancement in developing robust, adaptive investment strategies without human intervention.
The proposed framework demonstrates versatility across various asset classes, enhancing its utility and practical effectiveness. To support future research and ensure reproducibility, we make source code publicly available at https://github.com/kouzhizhuo/Automate-Strategy-Finding-with-LLM-in-Quant-investment.
2 Problem Formulation
This section establishes the theoretical foundation for our research on alpha factor strategies in quantitative finance. We formulate a framework addressing three interconnected challenges: mathematically formalizing these alphas, developing dynamic methodologies to generate seed alphas, and defining and optimizing alpha factor strategies. Our approach integrates LLMs and multi-agent systems to overcome limitations in traditional quantitative trading methods.
We employ consistent notation throughout: $n$ stocks observed over trading periods $t\in{1,2,\ldots,T}$, each characterized by $m$ financial features. Alpha factors are denoted as $\alpha_{ij}^{(t)}$ for stock $i$ in category $j$ at time $t$, with corresponding weights $w_{i}$. Market conditions at time $t$ are represented by $\mathcal{M}^{(t)}$, and alpha factor predictive power is measured using the Information Coefficient (IC) 1.
$IC=\sigma(u,v)$ (1)
A higher IC indicates stronger predictive relationships between alpha values and future returns. where $\sigma(u,v)$ is correlation coefficient between predicted alphas $u$ and actual future returns $v$.
2.1 Alpha Factor Representation
The first challenge involves effectively representing alpha factors using mathematical expressions that capture meaningful financial signals Lo (2007). Given raw financial features $\mathbf{X}i^{(t)}=\{X{i1}^{(t)},X_{i2}^{(t)},\ldots,X_{im}^{(t)}\}$ for each stock $i$ at time $t$ (such as price, volume, and volatility), we define alpha factors through two classes of operators:
Cross-Section Operators $f_{\text{cs}}$: These operators process data from a single time period, capturing instantaneous relationships between financial variables:
$\alpha_{cs}^{(t)}=f_{\text{cs}}(\mathbf{X}i^{(t)})$ (2)
Time-Series Operators $f_{\text{ts}}$: These operators analyze data spanning multiple periods, identifying trends and temporal patterns:
$\alpha_{ts}^{(t)}=f_{\text{ts}}(\mathbf{X}i^{(t)},\mathbf{X}_{i}^{(t-1)},\ldots,\mathbf{X}_{i}^{(t-n)})$ (3)
The seed alpha for stock $i$ in category $j$ at time $t$ is formulated as:
$\alpha_{ij}^{(t)}=w_{cs}\cdot f_{\text{cs}}(\mathbf{X}i^{(t)})+w{ts}\cdot f_{\text{ts}}(\mathbf{X}i^{(t)},\ldots,\mathbf{X}_{i}^{(t-n)})$ (4)
where $w_{cs}$ and $w_{ts}$ are weights assigned to the cross-section and time-series components respectively. This representation allows for the flexible combination of different financial signals, enabling the creation of complex and nuanced alpha factors.
2.2 Alpha Mining and Selection
The second challenge addresses the limitations of traditional alpha mining methods, which often fail to adapt to changing market conditions. We formulate a dynamic selection approach that identifies the most relevant alphas for current market conditions using two complementary evaluation mechanisms:
Confidence Score Evaluation: Assesses the statistical reliability of each alpha factor:
$\theta_{ij}=\mathbb{E}\left[IC(\alpha_{ij}^{(t)}|\mathcal{M}^{(t)})\right]$ (5)
where $\theta_{ij}$ represents the confidence score for alpha $\alpha_{ij}$, and $\mathbb{E}[\cdot]$ denotes the expected value. Higher confidence scores indicate more consistent performance across various market environments.
Risk Preference Evaluation: Examines the risk characteristics of each alpha factor:
$\rho_{ij}=f_{\text{risk}}(\alpha_{ij}^{(t)},\mathcal{M}^{(t)})$ (6)
where $\rho_{ij}$ represents the risk score and $f_{\text{risk}}$ is a function that evaluates how well the alpha performs under different risk scenarios.
The optimal set of seed alphas is selected by considering both confidence and risk evaluations:
$\alpha_{ij}=\underset{\alpha{ij}}{\arg\max}\,\left[w_{c}\cdot\theta_{ij}+w_{r}\cdot\rho_{ij}\right]$ (7)
where $w_{c}$ and $w_{r}$ are weights assigned to confidence and risk scores respectively, reflecting their relative importance in the selection process.
2.3 Strategy Optimization
The third challenge involves combining the selected alphas into an effective investment strategy. The final alpha strategy is defined as a weighted combination of the optimal alphas from each category:
$\alpha^{(t)}=\sum_{j=1}^{k}w_{j}\cdot\alpha_{ij}$ (8)
where $w_{j}$ represents the weight assigned to category $j$, and $k$ is the total number of alpha categories. These weights are dynamically adjusted based on current market status to maximize strategy performance while managing overall portfolio risk. This optimization completes our framework, transforming raw financial data into a robust trading strategy that can adapt to changing market environments.
Figure 1: Overview of the strategy generate process in three components, a Seed Alpha Factory built using LLMs, a multi-agent decision-making system, and a weight optimization approach for overall strategy (CS stands for confidence score; RP stands for risk preference)
3 Methodology
3.1 Framework Overview
Our framework comprises three interconnected components (Figure 1): the SAF, multi-agent decision-making, and weight optimization. The initial phase employs LLMs to analyze and categorize multimodal financial research documents, constructing a comprehensive SAF. The LLM’s capability to process diverse datasets enables the creation of a robust set of seed alphas categorized into independent groups, aligning with established finance alpha mining principles OpenAI (2023). The second phase implements a multimodal multi-agent evaluation process that incorporates varied risk perspectives, enhancing strategy adaptability across different market conditions. This phase produces an optimized alpha set tailored to current market states and risk preferences. The final phase employs deep learning methods to optimize the weights of selected alphas, constructing a cohesive overall strategy.
The framework’s dynamic architecture enables continuous refinement through incremental updates to the SAF as new research emerges and market conditions evolve. This adaptability ensures the strategy maintains relevance and robustness over time. The methodology’s versatility permits application to any structured market globally, effectively replicating and enhancing professional investment research approaches.
3.2 LLM-Based Seed Alpha Generation
The first stage implements an LLM-based filtering and categorization process for alpha-related research. We utilize GPT-4o model to perform these tasks on a diverse corpus of financial research. Our initial dataset comprises 11 documents spanning both theoretical and applied aspects of alpha mining research (see Appendix A.2). Through this process, the system generates 9 distinct categories containing 100 seed alphas. The approach enhances the model’s ability to extract intricate details and relationships within financial research, resulting in a more robust and diverse SAF.
The output is a structured set of seed alphas categorized into distinct financial domains such as Momentum, Mean Reversion, and Fundamental analysis et al. Each category includes specific alpha designations and corresponding executable formulations derived from the LLM’s analysis (see Appendix A.3). This structured output forms the foundation for subsequent processing stages.
3.3 Multimodal Multi-Agent Evaluation
The second stage implements a comprehensive evaluation and selection of alpha factors through a multimodal and multi-agent system. Our system incorporates five types of multimodal data (detailed in Appendix A.4) encompassing textual, numerical, visual, audio, and video inputs. The multi-agent architecture comprises two agents Confidence Score Agent (CSA) and Risk Preference Agent (RPA). Selected alpha factors undergo rigorous backtesting using historical market data, with key evaluation metrics including the IC Goodwin (1998) 1 and Sharpe Ratio Sharpe (1994) 9:
$\text{Sharpe Ratio}=\frac{\mathbb{E}[R(\alpha^{(t)})-R_{f}]}{\sqrt{\text{Var}[R(\alpha^{(t)})]}}$ (9)
where $R(\alpha^{(t)})$ represents the return of the strategy and $R_{f}$ represents the risk-free rate. We developed a Category-Based Alpha Selection algorithm (detailed in Appendix A.5) to automate the selection. This algorithm systematically identifies and selects alphas from different categories based on their confidence scores, ensuring rigorous selection of factors that meet confidence thresholds across all categories while maintaining category diversity.
3.4 Weight Optimization
The final stage employs a 3-layer MLP to optimize the weights of selected seed alphas by mapping historical alpha calculations to future yields Chen and et al. (2020). The architecture consists of an input layer processing daily alpha values, a hidden layer with ten ReLU-activated nodes, and a single-node output layer for yield prediction.
During training, the network employs backpropagation and gradient descent to minimize the loss function, quantifying the discrepancy between predicted and actual yields. We utilize a separate validation set to ensure model generalizability and prevent overfitting. The DNN processes input data through the hidden layer, transforming it with learned weights and biases, and generates the final output through the output layer’s weights, biases, and activation function.
This methodology establishes a robust framework for predicting future yields based on historical alpha values, demonstrating the efficacy of deep learning techniques in optimizing alpha weights and enhancing investment strategy performance. Algorithm LABEL:overallago provides a formal specification of our complete framework, illustrating the logical flow from multimodal data processing through multi-agent evaluation to final weight optimization.
4 Experiment
Our research aims to develop a comprehensive LLM-driven alpha mining framework that operates without human intervention. This framework is uniquely capable of processing multimodal information and adapting to varying market conditions. To validate the effectiveness of our framework, we have conducted a series of experiments.
4.1 Datasets
Our study focuses on financial data from the Chinese market and US market, specifically targeting the SSE 50 Index. Table 1 shows the experiment dataset, which encompasses six primary features as original inputs for our Alpha factors: open, high, low, close, volume (OHLCV), and volume-weighted average price (VWAP). To ensure rigorous evaluation and robust model performance. Our experiments integrate financial reports and factor performances of the 50 constituent companies of the SSE 50 Index, providing a comprehensive view of the market Li and Mei (2020). The evaluation considers various datasets, including financial reports from the specified periods and performance metrics of different Alpha factors.
Table 1: Summary of the experiment dataset
Aspect Details
Primary Features Open, High, Low, Close, Volume, VWAP
Alpha Factors Custom factors based on price, volume, financial ratios,
moving averages, sentiment analysis
Financial Reports Quarterly and Annual reports from Index constituent companies
Time Periods Jan 2019-Jun 2024
Market Coverage SSE50, CSI300, SP500 Index
Evaluation Criteria Causal relationships, Alpha factor performance, model robustness
4.2 Multimodal Knowledge Extraction and Adaptive Alpha Discovery
We implement a prompt architecture (Figure 2) that incorporates multimodal market information into LLMs for comprehensive knowledge extraction and seed alpha selection. This architecture integrates textual financial sentiment data, numerical company financial statements, and visual trading charts to provide holistic market analysis Luo et al. (2025b). Our contextual analysis mechanism dynamically adjusts parameters based on prevailing market trends and sector performance. Experimental validation demonstrates the framework’s adaptability across varying market conditions. In Case 1, analyzing SSE50 data from Jan 2023 to Dec 2023, the model selected momentum and volume-based indicators such as price momentum, RSI, and MACD. Conversely, in Case 2, when processing the time window from Jan 2022 to Dec 2022, the model prioritized volatility and economic factors, including ATR, Bollinger Bands, and gross profit indicators.
These results confirm the framework’s capacity to dynamically adapt to changing market conditions through effective multimodal data integration. This adaptability enables the identification of market-appropriate alphas, enhancing strategy robustness across diverse market environments.
Figure 2: Sample Experiment on Different Market Status Input and Alpha Selection, Selected Alpha Depends on Different Context
4.3 Comparative Performance Against Traditional Alpha Factories
When evaluating the performance of selected seed alpha signals, the primary metric is the IC. We evaluated five most common alpha categories: Momentum, Mean Reversion, Volatility, Fundamental and Growth. The results in Table 2 demonstrate that our LLM-driven framework consistently achieves higher average IC values across all categories, particularly in Volatility and Fundamental, indicating superior trading effectiveness compared to original.
Table 2: IC comparison of mean & selected
Momentum
Mean
Reversion
Volatility Fundamental Growth
Mean IC of SAF 0.0092 0.0135 0.0177 0.0118 0.0146
Mean IC of Selected SAF 0.0208 0.0187 0.0258 0.0192 0.0217
4.4 Performance Evaluation of the Integrated Alpha Framework in Benchmark Comparison
Figure 3: Cumulative return backtest result on SSE50. The line track the net worth of different methods
Table 3: SSE50 2023 test combination of 12 Alphas
#\# Alpha Weight IC(SSE50)
1 (CLOSE - DELAY(CLOSE, 14)) -0.1459 0.0209
2 (RSI - DELAY(RSI, 14)) -1.0265 -0.0225
3 (CLOSE - DELAY(SMA(CLOSE, 14), 7)) -0.1978 0.0193
4 (MA(CLOSE, 20) - CLOSE) 0.0556 -0.0186
5 (SMA(CLOSE, 20) - CLOSE) -0.945 -0.0186
6 (MAX(HIGH, 20) - CLOSE) -0.4053 -0.0185
7 (100-RSI) -0.3199 0.0194
8 (BOLL_UP - BOLL_DOWN) / SMA(CLOSE, 20) 3.6186 0.0278
9 STD(CLOSE, 10) / STD(CLOSE, 50) -0.183 0.0236
10 VOLUME / MARKET_CAP -3.2145 -0.0194
11 VOLUME * CLOSE -0.0058 0.0187
12 (EPS / DELAY(EPS, 1) - 1) -1.8351 -0.0215
Weighted Combination -0.0587
Table 3 presents an example combination of 12 alphas generated by our framework, evaluated on the SSE50 constituent stock set. The table details the seed alphas selected by the LLM from each category, along with their respective weights and IC 1 values. The weight combination IC 1 value is quite high as -0.0587. Although some of the seed alphas exhibit relatively low IC 1 values individually, their removal results in a significant drop in the retrained combination weight, indicating their critical role in the overall performance. For example, if we remove alpha #6 the weight combination will drop to -0.055; once we remove the alpha #11, the weight combination will drop to merely 0.0491. This suggests that the seed alpha set selected by the LLM synergizes effectively, providing robust predictive power Zhang et al. (2020). To address the question of whether our overall strategy, incorporating the LLM-driven framework and confidence scoring, can consistently beat the market, we conducted a backtest using a straightforward investment approach during the period from Jan 2023 to Jan 2024 (training: Jan 2021-Jun 2022; validation: Jul 2022-Dec 2022), on the SSE50 dataset we conduct the framework to construct the portfolio ( A.8). The net worth progression of the respective strategies over the testing period is shown in Figure 3, and the comparison matrix shows in Table 4. Despite not explicitly optimizing for absolute returns, our framework demonstrates impressive performance in the backtest, achieving the highest profit compared to other methods. Our cumulative return for 2023 backtest comes to $53.17\%$ positive and meanwhile the index performance is -11.73%. We also do the compression with 4 advanced statistical methods and 2 LLM based methods. The framework strategy substantially outperforms all competitors with the highest return (53.17%), best risk-adjusted metrics (Sharpe: 0.287, Sortino  [^2] Sortino ratio computing excess return per unit of downside deviation, evaluating only negative return volatility.: 0.208, Calmar  [^3] Calmar ratio divides annualized return by maximum drawdown, measuring reward relative to risk.: 1.052), and lowest volatility (0.762%). LLM based mathods FinCon and SEP show moderate success (22.47%, 17.89%), while conventional machine learning methodologies yield marginal returns (XGBoost 9.53%). Notably, all evaluated strategies surpass the SSE50 benchmark, which exhibits negative performance (-13.22%). This approach demonstrates the efficacy of our LLM-driven framework consistently outperforming market benchmarks through dynamic adaptation to changing conditions.
Table 4: Performance comparison of trading strategies
Strategy
Final
Return (%)
Sharpe
Ratio
Volatility
Sortino
Ratio
Calmar
Ratio
Ours 53.173 0.287 0.762 0.208 1.052
XGBoost Chen and Guestrin (2016) 9.532 0.038 1.019 0.067 0.103
LightGBM Ke et al. (2017) 7.125 0.030 0.993 0.053 0.066
MLP 3.110 0.013 0.960 0.023 0.043
PPO_filter Schulman et al. (2017) 2.865 0.013 0.886 0.024 0.017
FinCon Yu et al. (2024) 22.474 0.077 1.196 0.126 0.232
SEP Koa et al. (2024) 17.891 0.060 1.217 0.103 0.157
SSE 50 -13.22 -0.063 0.859 -0.111 -0.043
4.5 Framework Robustness Across Time Periods and Markets
We conducted rigorous cross-temporal and cross-market validation using two comprehensive datasets: CSI300  [^4] CSI300 Index tracking the 300 largest companies listed on China’s Shanghai and Shenzhen exchanges. constituents (accessed via Tushare API) and SP500  [^5] SP500 Index measuring the performance of 500 large U.S. companies traded on American stock exchanges. constituents (accessed via CRSP platform) spanning January 2019 to December 2023. Our dataset incorporated daily OHLCV price metrics, quarterly financial statements, and relevant macroeconomic indicators to enable robust alpha factor construction. To ensure methodological integrity, we systematically partitioned the dataset into training, validation, and testing subsets with precise chronological segmentation as detailed in Table 5. This experimental design facilitates comprehensive evaluation of our framework’s generalizability across different market environments and temporal contexts.
Table 5: Training, validation, and test Periods for CSI300 and SP500
Assets Training Period Validation Period Test Period
CSI300 Jan 2019-Jun 2020 Jun-Dec 2020 Jan-Jun 2021
Jan 2020-Jun 2021 Jun-Dec 2021 Jan-Jun 2022
Jan 2021-Jun 2022 Jun-Dec 2022 Jan-Jun 2023
SP500 Jan 2019-Jun 2020 Jun-Dec 2020 Jan-Jun 2021
Jan 2020-Jun 2021 Jun-Dec 2021 Jan-Jun 2022
Jan 2021-Jun 2022 Jun-Dec 2022 Jan-Jun 2023
We employed consistent temporal partitioning across Chinese A-share and US markets to enable direct comparative analysis. Our findings demonstrate the framework’s robust performance across diverse market environments. From Table 6 shows in the Chinese A-Share market, the strategy generated substantial alpha with CSI300 constituents, particularly in H1 2023 (192.27% annual return vs. benchmark 9.13%). Similarly, in the US market, the framework achieved strong returns with SP500 constituents (93.61% in H1 2021, 118.24% in H1 2023). Notably, during the H1 2022 market downturn, the strategy maintained positive returns in both markets (12.78% in CSI300, 2.77% in SP500) while their respective benchmarks declined significantly (-30.37% and -44.22%).
Table 6: Backtest performance results across different time windows
Time Window Annual Return (%) Cum. Return (%) Max DD (%)
Strategy Baseline Strategy Baseline Strategy Baseline
Our Strategy vs CSI300
Jan-Jun 2021 29.70 7.31 11.91 3.10 -19.40 -10.86
Jan-Jun 2022 12.78 -30.37 5.29 -14.37 -24.01 -28.59
Jan-Jun 2023 192.27 9.13 59.03 3.85 -17.03 -8.44
Our Strategy vs SP500
Jan-Jun 2021 93.61 29.67 35.19 12.59 -7.89 -4.23
Jan-Jun 2022 2.77 -44.22 1.25 -23.39 -20.55 -23.51
Jan-Jun 2023 118.24 35.22 42.78 14.76 -11.52 -7.75
Note:Max DD = Maximum Drawdown.
The empirical evidence substantiates both the temporal persistence and cross-market robustness of our findings, with notably beneficial countercyclical characteristics emerging during periods of market distress. The framework demonstrates consistent alpha generation across diverse market architectures, regulatory frameworks, and investor behavioral patterns, establishing its broad applicability. Specifically, the model maintains stable outperformance trajectories through bullish, bearish, and range-bound market conditions, thereby validating its structural integrity and adaptability to varying macroeconomic environments.
4.6 Ablation Study
Our ablation study systematically evaluates the contribution of CSA and RPA within a multi-agent framework. We evaluate three configurations of each, using the SSE50 index data (2010-2022) and measure performance by IC 1 and Sharpe Ratio 9.
Table 7 demonstrates that the complete model achieved superior performance, with an out-of-sample IC of 0.047 and Sharpe Ratio of 1.73. Removing the CSA caused substantial degradation, reducing out-of-sample IC by 31.9% and Sharpe Ratio by 22.5%. RPA removal also decreased performance metrics, with particularly significant impact on the Sharpe Ratio. These results indicate that while both components contribute meaningfully, the CSA plays a more critical role in maintaining predictive stability. To assess performance consistency across market conditions, we analyzed out-of-sample IC values during different market regimes (Table 8). The complete model maintained consistent performance across bull, bear, and sideways markets. In contrast, the model without CSA performed particularly poorly during bear markets(IC: 0.021 vs. 0.042). The model without RPA showed moderate degradation during non-bull markets, less severe than observed without CSA.
This analysis confirms that both components enhance performance stability, with the CSA providing especially critical functionality during adverse market conditions.
Table 7: Ablation study results: impact of agent components on performance metrics
Model Configuration
IC
(In-Sample)
IC
(Out-of-Sample)
Sharpe
Ratio
Full Model 0.059 0.047 1.94
Without Confidence Score 0.054 0.032 1.51
Without Risk Preference 0.056 0.039 1.34
Table 8: Performance across market regimes (out-of-sample IC)
Model Configuration
Bull
Market
Bear
Market
Sideways
Market
Full Model 0.051 0.042 0.045
Without Confidence Score 0.046 0.021 0.029
Without Risk Preference 0.049 0.028 0.037
4.7 Sensitive Study
We analyzed the framework’s sensitivity to agent weights and neural network hyperparameters. The CSA & RPA weight ratio of 0.6/0.4 provided superior performance across market regimes, particularly during bear markets (Sharpe: 10.37), indicating enhanced robustness to regime shifts (Table 9). Optimal neural network configuration (10 hidden nodes, learning rate: 0.001, batch size: 32, regularization: 0.001) achieved a Sharpe ratio of 13.33, with performance particularly sensitive to regularization strength (Table 10).
Table 9: Sensitivity to agent weight configurations
Confidence Risk Sharpe Ratio
Weight Weight Bull Market Bear Market Overall
1.0 0.0 4.32 6.60 8.10
0.8 0.2 -3.41 -7.23 -2.68
0.6 0.4 8.70 10.37 11.39
0.5 0.5 -0.49 -1.62 -5.10
0.4 0.6 -4.27 -4.41 -8.04
0.2 0.8 5.68 8.51 9.00
0.0 1.0 -0.61 1.78 -4.35
Table 10: Sensitivity to neural network hyperparameters
Hidden Learning Batch Regularization Sharpe
Nodes Rate Size Parameter Ratio
5 0.001 32 0.001 9.69
10 0.001 32 0.001 13.33
20 0.001 32 0.001 6.93
10 0.0005 32 0.001 7.03
10 0.002 32 0.001 5.13
10 0.001 16 0.001 6.26
10 0.001 64 0.001 4.25
10 0.001 32 0.0005 -1.88
10 0.001 32 0.002 -6.91
The analysis revealed regime-dependent behavior across market conditions, and critical sensitivity to regularization strength. The optimal configuration achieved a robust overall Sharpe ratio of 11.39.
5 Related Work
Formulaic alphas in quantitative investment represent systematic, rule-based strategies that generate excess returns by exploiting specific market patterns and inefficiencies Kakushadze (2016). These strategies employ various methodologies, including genetic programming that involves structural and numerical mutations to generate novel alphas Cong et al. (2021), enhanced time-series operators with mutual information as fitness measures Lin et al. (2019), and algorithmic graphs for more complex predictions Cui et al. (2021). Machine learning approaches utilize neural network architectures such as LSTM Hochreiter and Schmidhuber (1997) and Transformer models Vaswani et al. (2017), while decision tree models like XGBoost Chen and Guestrin (2016) and LightGBM Ke et al. (2017) offer interpretability advantages. Recent research focuses on integrating non-standard data sources, as demonstrated by REST Xu et al. (2021b) and HIST Xu et al. (2021a).
The development of general-domain LLMs has catalyzed interest in Finance LLMs (Fin-LLMs), although this specialized domain remains nascent Novy-Marx (2015); Yang et al. (2023); Zhao et al. (2023). Open-source LLMs such as LLaMA Touvron et al. (2023), BLOOM W. and et al. (2023), and Flan-T5 W. and et al. (2022) provide flexibility but may underperform proprietary alternatives. Fine-tuned financial LLMs demonstrate enhanced domain-specific comprehension, yet their generative performance indicates the need for improved domain-specific datasets Lewis and et al. (2020); Koa et al. (2024); Wen et al. (2025).
Multimodal LLMs have shown significant potential in investment contexts by processing diverse data types L et al. (2023); Luo et al. (2025a), developing strategies that mitigate market volatility K et al. (2024), and analyzing textual data to gauge investor sentiment Zhao et al. (2024). Multi-agent LLM systems enhance market analysis capabilities by leveraging vast datasets to interpret financial reports and market sentiment Zhang et al. (2024b), simulating various market scenarios Talebirad and Nadiri (2023), and facilitating parallel testing of diverse strategies Wang et al. (2024). Implementation raises important considerations regarding transparency, accountability, and bias mitigation Yu and et al. (2024); Mundhenk et al. (2021).
6 Conclusion
In this paper we proposed a novel quantitative investment framework integrating LLMs and multi-Agent architectures to address instability in traditional approaches. Our system generates diversified alpha factors from multimodal financial data, constructs risk-calibrated trading agents, and employs a deep learning mechanism for dynamic agent weighting based on market conditions.
Experimental results confirm the framework’s effectiveness across CN and US markets, our framework demonstrates significant outperformance versus SOTA alpha generation methods and benchmark indices across key financial metrics. This work successfully extends LLM capabilities to quantitative trading, creating a scalable, adaptive architecture for financial signal extraction that functions effectively without human intervention.
7 Limitations
Our framework presents several significant limitations. First, system efficacy is contingent upon input document quality, potentially perpetuating inherent biases Deb et al. (2017); Ashok et al. (2018). Second, LLM-generated alphas occasionally lack the financial intuition characteristic of human analysts, resulting in theoretically sound but practically infeasible factors Tuarob et al. (2017). Third, our multi-agent evaluation methodology presupposes persistent historical relationships between market conditions and alpha performance, an assumption that may prove tenuous during market regime shifts. Finally, our validation efforts have primarily targeted equity markets, with cross-asset applicability requiring additional empirical investigation. Future research directions should address these constraints through exploring Mixture of Experts (MoE) architectures to improve learning efficiency Masoudnia and Ebrahimpour (2014), adaptive agent architectures, transfer learning methodologies Wang et al. (2023), related regulation requirements Han and Xi (2020), and computationally efficient implementations.
8 Acknowledgments
This paper is partially supported by grants from the HKUST Start-up Fund (R9911), Theme-based Research Scheme grant (T45-205/21-N), the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government, and the research funding under HKUST-DXM AI for Finance Joint Laboratory (DXM25EG01). Also supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China (NSFC) under Grant No. U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No. 2024A0505040012, the Hong Kong RGC GRF Project 16213620, RIF Project R6020-19, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Key Areas Special Project of Guangdong Provincial Universities 2024ZDZX1006, Guangdong Province Science and Technology Plan Project 2023A0505030011, Guangzhou municipality big data intelligence key lab, 2023A03J0012, Hong Kong ITC ITF grants MHX/078/21 and PRP/004/22FX, Zhujiang scholar program 2021JC02X170, Microsoft Research Asia Collaborative Research Grant, HKUST-Webank joint research lab and 2023 HKUST Shenzhen-Hong Kong Collaborative Innovation Institute Green Sustainability Special Fund, from Shui On Xintiandi and the InnoSpace GBA. The authors are grateful to anonymous reviewers for their efforts and insightful suggestions to improve this article.
Appendix A Appendix
A.1 Sample of Seed Alpha
The seed alpha representation depicted in Figure 4 illustrates a fundamental construct in quantitative finance for generating trading signals. This figure presents a comprehensive visualization of the Detrended Price Oscillator (DPO) formula, which compares the current closing price with a delayed simple moving average (SMA) of closing prices. Panel A shows the mathematical formulation of the seed alpha, expressed as CLOSE - DELAY(SMA(CLOSE, 14), 7), which captures price momentum by measuring deviations from historical trends. Panel B transforms this formula into an equivalent expression tree, demonstrating the hierarchical relationship between operators and operands, which facilitates algorithmic implementation and analysis. Panel C provides a practical illustration through a tabulated step-by-step computation of this alpha formula on a sample time series, showing how the signal evolves over multiple trading days. This representation exemplifies how complex financial indicators can be systematically decomposed, formalized, and applied to market data for quantitative trading strategies.
Figure 4: Seed Alpha Representation: (A) An example of the seed alpha formula. (B) Its equivalent expression tree. (C) Step-by-step computation of this seed alpha on an example time series.
Category Name Short Code
Momentum Price Momentum (CLOSE - DELAY(CLOSE, 14))
Volume Momentum (VOLUME - DELAY(VOLUME, 14))
RSI Momentum (RSI - DELAY(RSI, 14))
Rate of Change (ROC) ((CLOSE / DELAY(CLOSE, 14)) - 1)
MACD Momentum (MACD - DELAY(MACD, 14))
Momentum Oscillator ((CLOSE - DELAY(CLOSE, 14)) / DELAY(CLOSE, 14))
Chande Momentum Oscillator (CMO) (SUM(IF(CLOSE - DELAY(CLOSE, 1) > 0, CLOSE - DELAY(CLOSE, 1), 0), 14) - SUM(IF(CLOSE - DELAY(CLOSE, 1) << 0, DELAY(CLOSE, 1) - CLOSE, 0), 14)) / (SUM(IF(CLOSE - DELAY(CLOSE, 1) > 0, CLOSE - DELAY(CLOSE, 1), 0), 14) + SUM(IF(CLOSE - DELAY(CLOSE, 1) << 0, DELAY(CLOSE, 1) - CLOSE, 0), 14)) * 100
Stochastic Momentum Index (SMI) ((CLOSE - MIN(LOW, 14)) - (MAX(HIGH, 14) - CLOSE)) / (MAX(HIGH, 14) - MIN(LOW, 14))
ATR Momentum (ATR - DELAY(ATR, 14))
Detrended Price Oscillator (DPO) (CLOSE - DELAY(SMA(CLOSE, 14), 7))
Average Directional Index (ADX) Momentum (ADX - DELAY(ADX, 14))
Mean Reversion Mean Reversion (MEAN(CLOSE, 20) - CLOSE)
Z-Score Mean Reversion (CLOSE - MEAN(CLOSE, 20)) / STD(CLOSE, 20)
Bollinger Bands (CLOSE - LOWER_BAND) / (UPPER_BAND - LOWER_BAND)
Keltner Channel (CLOSE - LOWER_CHANNEL) / (UPPER_CHANNEL - LOWER_CHANNEL)
Moving Average Reversion (SMA(CLOSE, 20) - CLOSE)
Exponential Moving Average (EMA) Reversion (EMA(CLOSE, 20) - CLOSE)
Distance from High (MAX(HIGH, 20) - CLOSE)
Distance from Low (CLOSE - MIN(LOW, 20))
Relative Strength Index (RSI) Reversion (100 - RSI)
Percent B ((CLOSE - LOWER_BAND) / (UPPER_BAND - LOWER_BAND)) * 100
Volatility Standard Deviation STD(CLOSE, 20)
Average True Range (ATR) ATR(14)
Bollinger Band Width (UPPER_BAND - LOWER_BAND) / SMA(CLOSE, 20)
Historical Volatility STD(RETURNS, 20) * SQRT(252)
Volatility Ratio STD(CLOSE, 10) / STD(CLOSE, 50)
Chaikin Volatility (EMA(HIGH - LOW, 10) / DELAY(EMA(HIGH - LOW, 10), 10)) - 1
Garman-Klass Volatility SQRT(0.5 * LOG(HIGH / LOW)2 - (2 * LOG(2) - 1) * LOG(CLOSE / OPEN)2)
Parkinson Volatility SQRT((1 / (4 * N * LOG(2))) * SUM(LOG(HIGH / LOW)2, 20))
Yang-Zhang Volatility SQRT(VAR(LOG(CLOSE / OPEN)) + 0.5 * VAR(LOG(HIGH / OPEN) - LOG(LOW / OPEN)) + 0.25 * VAR(LOG(CLOSE / DELAY(OPEN, 1))))
Ulcer Index SQRT(MEAN(DRAWDOWN2, 14))
Fundamental Price-to-Earnings Ratio (P/E) (CLOSE / EPS)
Price-to-Book Ratio (P/B) (CLOSE / BOOK_VALUE)
Dividend Yield (DIVIDENDS / CLOSE)
Earnings Yield (EPS / CLOSE)
Sales-to-Price Ratio (SALES / CLOSE)
Cash Flow Yield (OPERATING_CASH_FLOW / CLOSE)
Liquidity Trading Volume VOLUME
Average Trading Volume MEAN(VOLUME, 20)
Volume Rate of Change (VROC) (VOLUME - DELAY(VOLUME, 14)) / DELAY(VOLUME, 14)
On-Balance Volume (OBV) SUM(VOLUME * SIGN(CLOSE - DELAY(CLOSE, 1)))
Liquidity Ratio VOLUME / MARKET_CAP
Turnover Rate VOLUME / SHARES_OUTSTANDING
Amihud Illiquidity Ratio ABS(RETURN) / VOLUME
High-Low Spread (HIGH - LOW) / CLOSE
Dollar Volume VOLUME * CLOSE
Debt-to-Equity Ratio (TOTAL_DEBT / TOTAL_EQUITY)
Return on Equity (ROE) (NET_INCOME / EQUITY)
Return on Assets (ROA) (NET_INCOME / TOTAL_ASSETS)
Gross Profit Margin (GROSS_PROFIT / REVENUE)
Price-to-Sales Ratio (P/S) (CLOSE / SALES)
Price-to-Cash Flow Ratio (P/CF) (CLOSE / OPERATING_CASH_FLOW)
Book-to-Market Ratio (B/M) (BOOK_VALUE / CLOSE)
Enterprise Value to EBITDA (EV/EBITDA) (ENTERPRISE_VALUE / EBITDA)
Bid-Ask Spread (ASK_PRICE - BID_PRICE) / MID_PRICE
High-Low Spread (HIGH - LOW) / CLOSE
Dollar Volume VOLUME * CLOSE
Quality Gross Profit Margin (GROSS_PROFIT / REVENUE)
Operating Profit Margin (OPERATING_INCOME / REVENUE)
Net Profit Margin (NET_INCOME / REVENUE)
Earnings Stability STD(EPS, 5) / MEAN(EPS, 5)
Debt to Equity Ratio (TOTAL_DEBT / TOTAL_EQUITY)
Interest Coverage Ratio (EBIT / INTEREST_EXPENSE)
Cash Conversion Cycle (DIO + DSO - DPO)
Asset Turnover Ratio (REVENUE / TOTAL_ASSETS)
Growth Earnings Growth Rate (EPS / DELAY(EPS, 1) - 1)
Revenue Growth Rate (REVENUE / DELAY(REVENUE, 1) - 1)
EBITDA Growth Rate (EBITDA / DELAY(EBITDA, 1) - 1)
Cash Flow Growth Rate (CASH_FLOW / DELAY(CASH_FLOW, 1) - 1)
Dividends Growth Rate (DIVIDENDS / DELAY(DIVIDENDS, 1) - 1)
Book Value Growth Rate (BOOK_VALUE / DELAY(BOOK_VALUE, 1) - 1)
Sales Growth Rate (SALES / DELAY(SALES, 1) - 1)
Asset Growth Rate (ASSETS / DELAY(ASSETS, 1) - 1)
Equity Growth Rate (EQUITY / DELAY(EQUITY, 1) - 1)
Retained Earnings Growth Rate (RETAINED_EARNINGS / DELAY(RETAINED_EARNINGS, 1) - 1)
Technical Moving Average (MA) SMA(CLOSE, 20)
Exponential Moving Average (EMA) EMA(CLOSE, 20)
Relative Strength Index (RSI) RSI(14)
Moving Average Convergence Divergence (MACD) (EMA(CLOSE, 12) - EMA(CLOSE, 26))
Bollinger Bands UPPER_BAND - LOWER_BAND
Stochastic Oscillator ((CLOSE - MIN(LOW, 14)) / (MAX(HIGH, 14) - MIN(LOW, 14))) * 100
Average True Range (ATR) ATR(14)
Commodity Channel Index (CCI) (TYPICAL_PRICE - SMA(TYPICAL_PRICE, 20)) / (0.015 * MEAN_DEV(TYPICAL_PRICE, 20))
Williams %R ((MAX(HIGH, 14) - CLOSE) / (MAX(HIGH, 14) - MIN(LOW, 14))) * -100
Macro Economics GDP Growth Rate GDP - DELAY(GDP, n)
Inflation Rate CPI - DELAY(CPI, n)
Unemployment Rate UNEMPLOYMENT_RATE - DELAY(UNEMPLOYMENT_RATE, n)
Interest Rate INTEREST_RATE - DELAY(INTEREST_RATE, n)
Industrial Production Index IPI - DELAY(IPI, n)
Retail Sales Growth RETAIL_SALES - DELAY(RETAIL_SALES, n)
Housing Starts Growth HOUSING_STARTS - DELAY(HOUSING_STARTS, n)
Consumer Confidence Index (CCI) CCI - DELAY(CCI, n)
Trade Balance EXPORTS - IMPORTS
Foreign Exchange Reserves FX_RESERVES - DELAY(FX_RESERVES, n)
A.2 Seed Alphas document lists for GPTs
The presented table offers a comprehensive overview of cutting-edge research pertinent to the development of a SAF in quantitative finance. These documents collectively represent the confluence of traditional financial methodologies with advanced computational techniques. Of particular significance is Kakushadze’s "101 Formulaic Alphas," which provides a foundational repository of trading signals that can be categorized according to traditional financial factors. Lopez de Prado’s work on "Causal Factor Investing" introduces scientific rigor to factor classification, ensuring independence between alpha categories. The integration of artificial intelligence is evident in multiple studies, including "Alpha-GPT" and "FinGPT," which leverage LLMs for alpha generation. Portfolio construction methodologies are addressed through reinforcement learning approaches in "AlphaPortfolio" and "Dynamic Graph-based Deep Reinforcement Learning." Performance enhancement strategies are explored in "Mastering Stock Markets with Efficient Mixture of Diversified Trading Experts," while market dependency analysis is covered in "Model-Free Implied Dependence and the Cross-Section of Returns." This literature collection provides quantitative researchers with the theoretical frameworks and methodological tools necessary to construct a Seed Alpha Factory that systematically generates, categorizes, and implements trading signals across independent financial categories, balancing traditional financial theory with contemporary computational advancements.
Sample prompt "Summarize the document information to help quantitative researchers build the Seed Alpha Factory according to traditional financial categories, ensuring that each category of seed alphas is independent."
A.3 Generate Seed Alpha factory
This taxonomy presents quantitative trading signals organized into eight categories: Momentum, Mean Reversion, Volatility, Fundamental, Liquidity, Quality, Growth, Technical, and Macroeconomic indicators. Each category targets distinct market phenomena with specific mathematical implementations. Momentum factors identify persistent trends, Mean Reversion signals detect market overreactions, Volatility metrics quantify price dispersion, and Fundamental factors evaluate company valuations. The framework also includes Liquidity measures of trading activity, Quality indicators of operational efficiency, Growth metrics of financial expansion, Technical indicators derived from price-volume patterns, and Macroeconomic signals reflecting broader economic conditions. The mathematical formulations provided enable researchers to implement diverse, uncorrelated alpha factors for robust quantitative trading strategies.
A.4 Multimodal Data types
Table 11 presents a taxonomy of multimodal data types essential for comprehensive quantitative finance research. The classification encompasses five categories: textual data (financial reports, news articles, social media discourse), numerical data (historical price series, returns, volatility metrics), visual data (charts and technical analysis patterns), audio data (financial broadcasts and market commentary), and video data (specialized financial news channels). This multimodal framework enables researchers to develop more robust predictive models by integrating complementary information channels, potentially identifying market inefficiencies that remain undetectable when analyzing isolated data types. By synthesizing diverse information formats, quantitative analysts can gain multidimensional insights into market dynamics, enhancing both analytical depth and predictive capability in financial modeling applications.
Table 11: Multimodal data types
Data Type Description Examples
Textual Data Financial reports, academic papers, news articles, and other textual documents. Trading forums’ sentiment analysis and stock predictions, company disclosures, financial statements, Sina Finance
Numerical Data Historical stock market data, financial metrics, and performance indicators. Returns, log returns, annualized returns, volatility
Visual Data Charts, graphs, and other visual representations of financial data. Kline charts, trading charts
Audio Data Financial news broadcasts. Financial morning news radio, stock review radio, market discussion radio
Video Data Financial news channels. CCTV Securities Information Channel, CCTV News Broadcast (news affecting China’s stock market)
A.5 Category-Based Alpha Selection
Algorithm 1 delineates a systematic methodology for alpha selection in quantitative investment strategies, employing a category-based approach to ensure diversification across multiple financial factors. The procedure operates on a structured set of alpha categories $\mathcal{C}$, each representing distinct market phenomena such as momentum, mean reversion, or volatility. For each category, the algorithm identifies superior alpha factors through the SelectBestAlphas function, which presumably evaluates historical performance metrics. The innovation lies in the subsequent dual-agent evaluation system: the RiskPreferenceAgent assesses each alpha’s risk characteristics, while the ConfidenceScoreAgent evaluates the statistical robustness of its historical performance. These complementary evaluations are synthesized using weight parameters $w_{r}$ and $w_{c}$ to compute a comprehensive Final Score. Only alphas exceeding a predefined confidence threshold $X$ are incorporated into the final selection set $\mathcal{A}$. This methodical approach ensures that the resulting alpha portfolio exhibits both category diversification and individual signal quality, potentially enhancing risk-adjusted returns while mitigating exposure to specific market regimes or factor deterioration.
Algorithm 1 Category-Based Alpha Selection
1:Categories $\mathcal{C}=\{C_{1},\ldots,C_{m}\}$, each $C_{i}$ containing a set of alphas; confidence threshold $X$
2:Initialize selected alphas $\mathcal{A}\leftarrow\emptyset$
3:for each category $C_{i}\in\mathcal{C}$ do
4:  $\mathcal{A}_{i}\leftarrow$ SelectBestAlphas($C_{i}$)
5:  for each $\alpha\in\mathcal{A}_{i}$ do
6:   $\text{risk\_score}\leftarrow$ RiskPreferenceAgent($\alpha$)
7:   $\text{confidence\_score}\leftarrow$ ConfidenceScoreAgent($\alpha$)
8:   $\text{final\_score}\leftarrow w_{r}\cdot\text{risk\_score}+w_{c}\cdot\text{confidence\_score}$
9:   if $\text{final\_score}>X$ then
10:     $\mathcal{A}\leftarrow\mathcal{A}\cup\{\alpha\}$
11:   end if
12:  end for
13:end for
14:return $\mathcal{A}$
A.6 Dynamic Alpha Strategy Construction
Algorithm 2 Overall Framework Algorithm
1:Multimodal financial data $\mathcal{D}$, market conditions $\mathcal{M}^{(t)}$, stocks $\mathcal{S}$, confidence threshold $\tau$
2:Optimized alpha strategy $\alpha^{(t)}$
3:/* Phase 1: SAF */
4:Initialize empty Seed Alpha Factory $\mathcal{F}\leftarrow\emptyset$
5:for each document $d\in\mathcal{D}$ do
6:  filtered_content $\leftarrow$ LLM.Filter($d$)
7:  categories $\leftarrow$ LLM.Categorize(filtered_content)
8:  for each category $c\in$ categories do
9:   seed_alphas $\leftarrow$ LLM.GenerateAlphas($c$)
10:   $\mathcal{F}_{c}\leftarrow\mathcal{F}_{c}\cup$ seed_alphas
11:  end for
12:end for
13:/* Phase 2: Multi-Agent Evaluation */
14:Initialize selected alphas $\mathcal{A}\leftarrow\emptyset$
15:for each category $c\in\mathcal{F}$ do
16:  for each alpha $\alpha_{ij}\in\mathcal{F}_{c}$ do
[…]
19:   $\text{score}_{ij}\leftarrow w_{c}\cdot\theta_{ij}+w_{r}\cdot\rho_{ij}$ $\triangleright$ Combined score
20:   if $\text{score}_{ij}>\tau$ then
21:     $\mathcal{A}_{c}\leftarrow\mathcal{A}_{c}\cup\{\alpha_{ij}\}$
22:   end if
23:  end for
24:  if $\mathcal{A}_{c}\neq\emptyset$ then
25:   $\alpha_{c}^{*}\leftarrow\arg\max\limits_{\alpha_{ij}\in\mathcal{A}_{c}}\text{score}_{ij}$ $\triangleright$ Select best alpha in category
26:   $\mathcal{A}\leftarrow\mathcal{A}\cup\{\alpha_{c}^{*}\}$
27:  end if
28:end for
29:/* Phase 3: Weight Optimization */
30:Initialize MLP with architecture $\{|\mathcal{A}|,10,1\}$ $\triangleright$ Input, hidden, output layers
[…]
34:for each stock $S_{i}\in\mathcal{S}$ do
35:  $\alpha_{i}^{(t)}\leftarrow\sum_{j=1}^{|\mathcal{A}|}w_{j}\cdot\alpha_{ij}^{(t)}$ $\triangleright$ Compute composite alpha
36:end for
37:return $\alpha^{(t)}=\{\alpha_{1}^{(t)},\alpha_{2}^{(t)},\ldots,\alpha_{n}^{(t)}\}$ $\triangleright$ Final alpha strategy
A.7 Sample prompt for Seed Alpha selection
1. The training set includes:
  • Financial Reports: 6 quarters (from Jan 2021 to Jun 2022) for 50 companies listed in the SSE 50.
  • Factor Analysis Data: 37 factors (from Jan 2021, to Jun 2022) divided into groups: Momentum, Mean Reversion, Volatility, Fundamental, Quality, Growth, Technical, Macro Economics. The metric used is the IC.
2. Objective: Learn the relationship between the performance of financial reports for the first four quarters from 2022 to 2023 and the factor analysis data (IC) for each of the 37 factors in the last quarter of 2022.
3. When provided with the test set (performance of the first 4 quarters of 2023):
  • Select the factors that will perform best in the last quarter of the SSE 50.
  • Provide a confidence score & the risk preference for your selection for each selected Alpha factor.
4. Selection Criteria:
  • If no relationship between financial reports and IC can be found, select the Alpha factor with the highest IC value in each group.
  • For verification of market information differences, if no relationship between financial reports and IR can be found, select the Alpha factor with the highest IR value.
A.8 Portfolio Construction Methods
Our investment methodology implements a daily portfolio reconstruction using a top-kk/drop-$n$ selection framework. At each trading session, we rank all securities by their alpha values—quantitative indicators of expected excess returns—and select the top $k$ stocks for portfolio inclusion. This approach targets securities with the strongest signals, potentially exploiting short-term market inefficiencies. We employ an equal-weighting scheme across selected securities, distributing capital homogeneously among the top candidates, which offers diversification benefits and aligns with our view that alpha signals primarily provide directional indications rather than precise return forecasts.
To optimize transaction costs and operational efficiency, we limit portfolio turnover to a maximum of $n$ securities per trading day. This constraint balances maintaining alignment with current alpha signals while minimizing trading friction costs. For our experiment in Section 4.4, we set $k=13$ and $n=5$ based on extensive backtesting. This configuration establishes a portfolio concentration that balances diversification against signal dilution, while limiting daily turnover to approximately 38%—a level that demonstrates favorable characteristics in our transaction cost modeling and represents an efficient trade-off between signal utilization and implementation costs.